OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (lncognlto):

First we need to substitute the info given into the formula y = mx + c, where m is slope and c is the y intercept. so 5 = 3(2) + c. Solving for c gives us c = 5 - 6 = -1. Now substituting c = -1 and m = 3 into the original formula, y = mx + c, what do we get?

OpenStudy (anonymous):

we get 5=5 right ?

OpenStudy (lncognlto):

Well, yes, we do, and this proves that the equation is correct, but I was thinking more along the lines of y = 3x - 1, which would be the equation of the line in point-slope form.

OpenStudy (anonymous):

y + 8 = 3(x + 5 right

OpenStudy (lncognlto):

Explain?